# Methods to integrate multinormals and compute classification measures

### v11

@@ Abstract
Univariate and multivariate normal probability distributions are widely used when modeling decisions under uncertainty. Computing the performance of such models requires integrating these distributions over specific domains, which can vary widely across models. Besides some special cases, there exist no general analytical expressions, standard numerical methods or software for these integrals. Here we present mathematical results and open-source software that provide (i) the probability in any domain of a normal in any dimensions with any parameters, (ii) the probability density, cumulative distribution, and inverse cumulative distribution of any function of a normal vector, (iii) the classification errors among any number of normal distributions, the Bayes-optimal discriminability index and relation to the operating characteristic, (iv) dimension reduction and visualizations for such problems, and (v) tests for how reliably these methods may be used on given data. We demonstrate these tools with vision research applications of detecting occluding objects in natural scenes, and detecting camouflage.
@@ 1 Introduction
The univariate or multivariate normal (henceforth called simply ‘normal’) is arguably the most important and widely-used probability distribution. It is frequently used because various central-limit theorems guarantee that normal distributions will occur commonly in nature, and because it is the simplest and most tractable distribution that allows arbitrary correlations between the variables.
Normal distributions form the basis of many theories and models in the natural and social sciences. For example, they are the foundation of Bayesian statistical decision/classification theories using Gaussian discriminant analysis [2], and are widely applied in diverse fields such as vision science, neuroscience, probabilistic planning in robotics, psychology, and economics. These theories specify optimal performance under uncertainty, and are often used to provide a benchmark against which to evaluate the performance (behavior) of humans, other animals, neural circuits or algorithms. They also serve as a starting point in developing other models/theories that describe sub-optimal performance of agents.
To compute the performance predicted by such theories it is necessary to integrate the normal distributions over specific domains. For example, a particularly common task in vision science is classification into two categories (e.g. detection and discrimination tasks). The predicted maximum accuracy in such tasks is determined by integrating normals over domains defined by a quadratic decision boundary [3, 4]. Predicted accuracy of some of the possible sub-optimal models is determined by integrating over other domains.
@@
Thus, there is no single analytical expression, numerical method or standard software tool to quickly and accurately integrate arbitrary normals over arbitrary domains, or to compute classification errors and the discriminability index $d\textquoteright$ ’. Evaluating these quantities is often simplified by making the limiting assumption of equal variance. This impedes the quick testing, comparison, and optimization of models. Here we describe a mathematical method and accompanying software implementation which provides functions to (i) integrate normals with arbitrary means and covariances in any number of dimensions over arbitrary domains, (ii) compute the pdf, cdf and inverse cdf of any function of a multinormal variable (normal vector), and (iii) compute the performance of classifying amongst any number of normals. This software is available as a Matlab toolbox ‘Integrate and classify normal distributions’, and the source code is at github.com/abhranildas/IntClassNorm.
We first review and assimilate previous mathematical results into a generalized chi-squared method that can integrate arbitrary normals over quadratic domains. Then we present our ray-trace method to integrate arbitrary normals over any domain, and consequently to compute the distribution of any real-valued function of a normal vector. We describe how these results can be used to compute error rates (and other relevant quantities) for Bayes-optimal and custom classifiers, given arbitrary priors and outcome cost matrix. We then present some methods to reduce problems to fewer dimensions, for analysis or visualization. Next, we provide a way to test whether directly measured samples from the actual distributions in a classification problem are close enough to normal to trust the computations from the toolbox. After describing the methods and software toolbox with examples, we demonstrate their accuracy and speed across a variety of problems. We show that for quadratic-domain problems both the generalized chi-squared method and the ray-trace method are accurate, but vary in relative speed depending on the particular problem. Of course, for domains that are not quadratic, only the ray-trace method applies. Finally, we illustrate the methods with two applications from our laboratory: modeling detection of occluding targets in natural scenes, and detecting camouflage.
@@ 2.1 In quadratic domains: the generalized chi-square method
The required probability, $p\left(\tilde{\chi}>0\right)$, is now a 1d integral, computable using, say, Ruben’s [8] or Davies’ [9] methods. We use the Matlab toolbox ‘Generalized chi-square distribution’ that we developed (source code is at github.com/abhranildas/gx2), which can compute the generalized chi-square parameters corresponding to a quadratic form of a normal vector, its statistics, cdf (using three different methods), pdf, inverse cdf and random numbers.
@@ 2.2.1 Standard polar form
where $\tilde{\Omega}$ is the domain where $\tilde{f}(\bm{z})>0$, and $\tilde{\Omega}_{\bm{n}}$ is its slice along the axis $\bm{n}$, i.e. the intervals along the axis where the axial domain function $\tilde{f}_{\bm{n}}(z)=\tilde{f}(z\bm{n})>0$. This may be called the ‘standard polar form’ of the integral. $d\bm{n}$ is the differential angle element ($d\theta$ in 2d, $\sin\theta\,d\theta\,d\phi$ in 3d etc).
@@ 2.2.3 Standard normal distribution on a ray
Consider the probability in the angular slice $d\bm{n}$ below some negative $z$ such as $z_{1}$ in fig. 1a. Note that the probability of a standard normal beyond some radius is given by the chi distribution. If $\Omega_{k}$ is the total angle in $k$ dimensions (2 in 1d, $2\pi$ in 2d, $4\pi$ in 3d), and $F_{\chi_{k}}(x)$ is the cdf of the chi distribution with $k$ degrees of freedom, we have:
@@ 2.2.4 Probability in an angular slice
Beyond 3d, we can use Monte Carlo integration over the angles. We draw a sample of random numbers from the standard multinormal in those dimensions, then normalize their magnitudes, to get a uniform random sample of rays $\bm{n}$, over which the expectation $\langle\alpha(\bm{n})\rangle/2$ is the probability estimate. Fig. 2b shows the computation of the 4d standard normal probability in the domain $f_{p}(\bm{x})=\sum_{i=1}^{4}\lvert x_{i}\rvert<1$, a 4d extension of a regular octahedron with plane faces meeting at sharp edges.
@@ 2.2.6 Probabilities of functions of a normal vector
We previously mentioned the equivalent ‘normal probability’ and ‘function probability’ views of conceptualizing a normal integral. So far we have mostly used the normal probability view, seeing scalar functions $f(\bm{x})$ as defining integral domains of the normal $\bm{x}$. But in the function probability view, $f(\bm{x})$ is instead seen as a mapping from the multi-dimensional variable $\bm{x}$ to a scalar, which can be considered a decision variable. Hence, integrating the normal in the multi-dimensional domain $f(\bm{x})>0$ corresponds to integrating the 1d pdf of the decision variable $f(\bm{x})$ beyond 0. It is helpful to plot this 1d pdf, especially when there are too many dimensions of $\bm{x}$ to visualize the normal probability view.
Conversely, given any scalar function $f(\bm{x})$ of a normal, its cdf, $F_{f}(c)=p(f(\bm{x})<c)$, is computed as the normal probability in the domain $c-f(\bm{x})>0$. Differentiating this gives us the pdf. (If it is a quadratic function, its generalized chi-square pdf can also be computed by convolving the constituent noncentral chi-square pdf’s.) Figs. 2a-c and g show 1d pdf’s of functions computed in this way. Also, inverting the function cdf using a numerical root-finding method gives us its inverse cdf.
@@ 4 Matlab toolbox: functions and examples
For an integration problem, the toolbox provides the function integrate_normal that inputs the normal parameters and the integration domain (as quadratic domain coefficients, limits of a rectangular domain, a ray-trace domain function or an implicit domain function), and outputs the integral and its complement, the boundary points computed, and a plot of the normal probability or function probability view. The toolbox also provides functions norm_fun_pdf, norm_fun_cdf and norm_fun_inv to compute pdf’s, cdf’s and inverse cdf’s of functions of (multi)normals. The function classify_normals for a binary classification problem inputs normal parameters, priors, outcome values and an optional classification boundary, and outputs the coefficients of the quadratic boundary $\beta$ and points on it, the error matrix and discriminability indices $d^{\prime}_{b}$, $d^{\prime}_{a}$ and $d^{\prime}_{e}$, and produces a normal probability or function probability plot. Optionally, it can also return the contributions to $d^{\prime}$ from each dimension along with a divided color-bar in the plot. With sample input, it first estimates the normal parameters (means, covariances and priors) and returns the optimal boundary $\beta$, and also which sample points were correctly classified. In addition, it returns the coefficients of the sample-optimal boundary $\gamma$ and points on it, error matrices and $d^{\prime}_{b}$ values of classifying the samples using $\beta$ and $\gamma$, and the mapped scalar decision variables $\beta(\bm{x})$ and $\gamma(\bm{x})$ from the samples. The function classify_normals_multi is similar and for classifying multiple normals.
Many different example problems, including every problem discussed in this paper (examples in figs. 2, 3, 4 and 5, tests in figs. 6 and 7, and research applications in fig. 8) are available as interactive demos in the ‘getting started’ live script of the toolbox, and can be easily adapted to other problems.
@@ 5 Performance benchmarks
In this section we test the performance (accuracy and speed) of our Matlab implementations of the generalized chi-square and ray-trace algorithms, against a standard Monte Carlo integration algorithm. We first set up a case where we know the ground truth, and compare the estimates of all three methods at the limit of high discriminability (low error rates) where it is most challenging (which occurs for computational models and ideal observers). We take two 3d normals with the same covariance matrix, so that the true discriminability $d^{\prime}$ is exactly calculated as their Mahalanobis distance. Now we increase their separation, while computing the optimal error with each of our methods at maximum precision, and a discriminability from it. The generalized chi-square method is very fast (due to the trivial planar boundary here), and the ray-trace method takes an average of 40s. For a fair comparison, we use $10^{8}$ samples for the Monte Carlo, which also takes $\sim$40s. Each method returns an estimate $\hat{d}^{\prime}$. Fig. 7a shows the relative inaccuracies $\lvert\hat{d}^{\prime}-d^{\prime}\rvert/d^{\prime}$ as true $d^{\prime}$ increases. With increasing separation, the Monte Carlo method quickly becomes inaccurate, since the error rate, i.e. the probability content in the integration domain, becomes too small to be sampled. The method stops working beyond $d^{\prime}\approx 10$, where none of the $10^{8}$ samples fall in the error domain. In contrast, inaccuracies in our methods are extremely small, of the order of the double-precision machine epsilon $\epsilon$, demonstrating that the algorithms contain no systematic error besides machine imprecision (however, Matlab’s native integration methods may not always reach the desired precision for a problem). This is possible because a variety of techniques are built into our algorithms to preserve accuracy, such as holding tiny and large summands separate to prevent rounding, using symbolic instead of numerical calculations, and using accurate tail probabilities. The inaccuracies do not grow with increasing separation, until $d^{\prime}\approx 75$, which corresponds to the smallest error rate $p(e)$ representable in double-precision ( realmin = 2e-308), beyond which both methods return $p(e)=0$ and $d^{\prime}_{b}=\infty$. For 1d problems beyond this, the toolbox returns $d^{\prime}_{e}$ instead.

### v12

@@ Abstract
Univariate and multivariate normal probability distributions are widely used when modeling decisions under uncertainty. Computing the performance of such models requires integrating these distributions over specific domains, which can vary widely across models. Besides some special cases, there exist no general analytical expressions, standard numerical methods or software for these integrals. Here we present mathematical results and open-source software that provide (i) the probability in any domain of a normal in any dimensions with any parameters, (ii) the probability density, cumulative distribution, and inverse cumulative distribution of any function of a normal vector, (iii) the classification errors among any number of normal distributions, the Bayes-optimal discriminability index and relation to the operating characteristic, (iv) ways to scale the discriminability of two distributions, (v) dimension reduction and visualizations for such problems, and (vi) tests for how reliably these methods may be used on given data. We demonstrate these tools with vision research applications of detecting occluding objects in natural scenes, and detecting camouflage.
@@ 1 Introduction
The univariate or multivariate normal (henceforth called simply ‘normal’) is arguably the most important and widely‐used probability distribution. It is frequently used because various central‐limit theorems guarantee that normal distributions will occur commonly in nature, and because it is the simplest and most tractable distribution that allows arbitrary correlations between the variables.
Normal distributions form the basis of many theories and models in the natural and social sciences. For example, they are the foundation of Bayesian statistical decision/classification theories using Gaussian discriminant analysis [2], and are widely applied in diverse fields such as vision science, neuroscience, probabilistic planning in robotics, psychology, and economics. These theories specify optimal performance under uncertainty, and are often used to provide a benchmark against which to evaluate the performance (behavior) of humans, other animals, neural circuits or algorithms. They also serve as a starting point in developing other models/theories that describe sub‐optimal performance of agents.
To compute the performance predicted by such theories it is necessary to integrate the normal distributions over specific domains. For example, a particularly common task in vision science is classification into two categories (e.g. detection and discrimination tasks). The predicted maximum accuracy in such tasks is determined by integrating normals over domains defined by a quadratic decision boundary [3, 4]. Predicted accuracy of some of the possible sub‐optimal models is determined by integrating over other domains.
@@
Thus, there is no single analytical expression, numerical method or standard software tool to quickly and accurately integrate arbitrary normals over arbitrary domains, or to compute classification errors and the discriminability index $d’$ ’. Evaluating these quantities is often simplified by making the limiting assumption of equal variance. This impedes the quick testing, comparison, and optimization of models. Here we describe a mathematical method and accompanying software implementation which provides functions to (i) integrate normals with arbitrary means and covariances in any number of dimensions over arbitrary domains, (ii) compute the pdf, cdf and inverse cdf of any function of a multinormal variable (normal vector), and (iii) compute the performance of classifying amongst any number of normals. This software is available as a Matlab toolbox ‘Integrate and classify normal distributions’, and the source code is at github.com/abhranildas/IntClassNorm.
We first review and assimilate previous mathematical results into a generalized chi-squared method that can integrate arbitrary normals over quadratic domains. Then we present our ray-trace method to integrate arbitrary normals over any domain, and consequently to compute the distribution of any real‐valued function of a normal vector. We describe how these results can be used to compute error rates (and other relevant quantities) for Bayes-optimal and custom classifiers, given arbitrary priors and outcome cost matrix. We then present some methods to reduce problems to fewer dimensions, for analysis or visualization. Next, we provide a way to test whether directly measured samples from the actual distributions in a classification problem are close enough to normal to trust the computations from the toolbox. After describing the methods and software toolbox with examples, we demonstrate their accuracy and speed across a variety of problems. We show that for quadratic-domain problems both the generalized chi-squared method and the ray-trace method are accurate, but vary in relative speed depending on the particular problem. Of course, for domains that are not quadratic, only the ray-trace method applies. Finally, we illustrate the methods with two applications from our laboratory: modeling detection of occluding targets in natural scenes, and detecting camouflage.
@@ 2.1 In quadratic domains: the generalized chi-square method
In a later paper [8], we derive the inverse map, i.e. from the parameters of a generalized chi-square distribution, to the parameters of the corresponding quadratic form of a multinormal.
The required probability, $p\left(\tilde{\chi}>0\right)$, is now a 1d integral, computable using, say, Ruben’s [9], Imhof’s [10, 11], IFFT [8] or ray [8] methods. We use the Matlab toolbox ‘Generalized chi-square distribution’ that we developed (source code is at github.com/abhranildas/gx2), which can compute the generalized chi-square parameters corresponding to a quadratic form of a normal vector, its statistics, cdf (using three different methods), pdf, inverse cdf and random numbers.
@@ 2.2.1 Standard polar form
where $\tilde{\Omega}$ is the domain where $\tilde{f}(\bm{z})>0$, and $\tilde{\Omega}_{\bm{n}}$ is its slice along the axis $\bm{n}$, i.e. the intervals along the axis where the axial domain function $\tilde{f}_{\bm{n}}(z)=\tilde{f}(z\bm{n})>0$. This may be called the ‘standard polar form’ of the integral. $d\bm{n}$ is the differential angle element ($d\theta$ in 2d, $\sin\theta\,d\theta\,d\phi$ in 3d, $\sin\theta\,\sin^{2}\psi\,d\theta\,d\phi\,d\psi$ in 4d etc).
@@ 2.2.3 Standard normal distribution on a ray
Consider the probability in the angular slice $d\bm{n}$ below some negative $z$ such as $z_{1}$ in fig. 1a. Note that the probability of a standard normal beyond some radius is given by the chi distribution. If $\Omega_{k}$ is the total angle in $k$ dimensions (2 in 1d, $2\pi$ in 2d, $4\pi$ in 3d, $2\pi^{2}$ in 4d), and $F_{\chi_{k}}(x)$ is the cdf of the chi distribution with $k$ degrees of freedom, we have:
@@ 2.2.4 Probability in an angular slice
Beyond 4d (or even in general), we can use Monte Carlo integration over the angles. We draw a sample of random numbers from the standard multinormal in those dimensions, then normalize their magnitudes, to get a uniform random sample of rays $\bm{n}$, over which the expectation $\langle\alpha(\bm{n})\rangle/2$ is the probability estimate. Fig. 2b shows the computation of the 4d standard normal probability in the domain $f_{p}(\bm{x})=\sum_{i=1}^{4}\lvert x_{i}\rvert<1$, a 4d extension of a regular octahedron with plane faces meeting at sharp edges. This Monte-Carlo integration can be done in parallel. So, just like ray-tracing for graphics rendering, this method can run on the GPU, and automatically uses one when available, batching the rays to prevent memory overflow. This provides up to a 4x speedup.
@@ 2.2.6 Probabilities of functions of a normal vector
We previously mentioned the equivalent ‘normal probability’ and ‘function probability’ views of conceptualizing a normal integral. So far we have mostly used the normal probability view, seeing scalar functions $f(\bm{x})$ as defining integral domains of the normal $\bm{x}$. But in the function probability view, $f(\bm{x})$ is instead seen as a mapping from the multi-dimensional variable $\bm{x}$ to a scalar, which in some contexts can be considered a decision variable. Hence, integrating the normal in the multi-dimensional domain $f(\bm{x})>0$ corresponds to integrating the 1d pdf of the decision variable $f(\bm{x})$ beyond 0. It is sometimes helpful to plot this 1d pdf, especially when there are too many dimensions of $\bm{x}$ to visualize the normal probability view.
Conversely, given any scalar function $f(\bm{x})$ of a normal, its cdf, $F_{f}(c)=p(f(\bm{x})<c)$, is computed as the normal probability in the domain $c-f(\bm{x})>0$. From this, we can obtain the pdf by numerically differentiating the cdf. In a later paper, we also analytically differentiate the ray-tracing method to directly compute the pdf of a scalar function of a normal vector. [8] (If it is a quadratic function, its generalized chi-square pdf can also be computed by convolving the constituent noncentral chi-square pdf’s.) Figs. 2a-c and g show 1d pdf’s of functions computed in this way. Also, inverting the function cdf using a numerical root-finding method gives us its inverse cdf.
@@ 3.2 Classifying using data
3.3 Scaling the discriminability of data
Figure 6: Scaling the discriminability of two distributions by linear interpolation. a. The discriminability of two general distributions (the orange one is non-normal) is scaled by linearly interpolating the mean and sd matrix of the orange towards the blue. Ellipses show their error ellipses. b. Scaling the discriminability of the data distributions in plot a by computing their decision variables under the normal model, and pulling them towards 0. $p(e)$ is the error rate.
Consider two samples $\bm{x}_{a}$ and $\bm{x}_{b}$, with means $\bm{\mu}_{a}$, $\bm{\mu}_{b}$ and covariances $\bm{\Sigma}_{a}$, $\bm{\Sigma}_{b}$ (sd matrices $\mathbf{S}_{a}$, $\mathbf{S}_{b}$). In some cases, we may want to scale their discriminability by moving them closer or farther apart. One such case is when we are modeling a detection or classification task, and the model performance exceeds that of the subject or observed data. In that case, we can move the model variable distributions closer together so that it matches the observed performance, while also predicting which specific data points should start overlapping and be misclassified.
To do this, we first model the samples as multinormals: $\bm{x}_{a}\sim\mathbf{S}_{a}\bm{z}+\bm{\mu}_{a}$ and $\bm{x}_{b}\sim\mathbf{S}_{b}\bm{z}+\bm{\mu}_{b}$. Now we effect transformations to linearly interpolate these multinormals closer together. There are different flavours of doing this depending on the case. If we suppose that $a$ is the noise distribution, then it should stay fixed while the signal distribution $b$ is pulled towards it. So we introduce the interpolation factor $r$, such that the parameters of distribution $b$ become: $\bm{\mu}^{\prime}_{b}=r\bm{\mu}_{b}+(1-r)\bm{\mu}_{a}$ and $\mathbf{S}^{\prime}_{b}=r\mathbf{S}_{b}+(1-r)\mathbf{S}_{a}$. As $r$ goes to 0, distribution $b$ linearly morphs into $a$. We can effect this transformation on the data points $\bm{x}_{b}$ by first whitening them: $\bm{z}_{b}=\mathbf{S}_{b}^{-1}(\bm{x}_{b}-\bm{\mu}_{b})$, then linearly transforming them to the new parameters: $\bm{x}^{\prime}_{b}=\mathbf{S}^{\prime}_{b}\bm{z}_{b}+\bm{\mu}^{\prime}_{b}$. We can also have other flavours, such as interpolating the parameters of the two distributions towards their common means. One important thing to note here is that the distribution parameters that we should linearly interpolate are the mean vector and the sd matrix, not the covariance matrix, because the points are a linear function of these parameters.
Fig. 6a shows two distributions where this linear warping is implemented. These can be general distributions. Here, for example, the blue one is normal while the orange one isn’t. The orange distribution is interpolated towards the blue in steps from $r=1$ through 0.05, and the $d^{\prime}$ shrinks. Using the optimal classification boundary at any step in the scaling will tell us which specific data points would be misclassified in the model, which we can then compare with our subject or observed data.
A different way that we can scale the discriminability of two data distributions in one dimension instead of many, is by computing their decision variables (log likelihood ratio $l$ that a point belongs to $a$ vs $b$) under the normal model, then moving them closer together or farther apart. Fig. 6b, top subplot, shows the two distributions of the Bayes decision variable for the two original bivariate samples in plot a that were farthest apart. (The blue sample $a$ is now on the right because its log likelihood ratio is positive.) If the median value of a decision variable distribution is $l_{m}$, we can map the decision variables to $l^{\prime}=l-l_{m}(1-r)$. As the scale factor $r$ goes from 1 to 0, each decision variable distribution is pulled in until half of its points are on either side of 0, i.e. the error rate goes to chance and the discriminability goes to 0. Once again, we can use this to scale down the discriminability of a model to match a subject or observed data, and see if the model predicts the specific trials or data points that were misclassified.
@@ 4 Matlab toolbox: functions and examples
For an integration problem, the toolbox provides the function integrate_normal that inputs the normal parameters and the integration domain (as quadratic domain coefficients, limits of a rectangular domain, a ray-trace domain function or an implicit domain function), and outputs the integral and its complement, the boundary points computed, and a plot of the normal probability or function probability view. The toolbox also provides functions norm_fun_pdf, norm_fun_cdf and norm_fun_inv to compute pdf’s, cdf’s and inverse cdf’s of functions of (multi)normals. The function classify_normals for a binary classification problem inputs normal parameters, priors, outcome values and an optional classification boundary, and outputs the coefficients of the quadratic boundary $\beta$ and points on it, the error matrix and discriminability indices $d^{\prime}_{b}$, $d^{\prime}_{a}$ and $d^{\prime}_{e}$, and produces a normal probability or function probability plot. Optionally, it can also return the contributions to $d^{\prime}$ from each dimension along with a divided color-bar in the plot. With sample input, it first estimates the normal parameters (means, covariances and priors) and returns the optimal boundary $\beta$, and also which sample points were correctly classified. In addition, it returns the coefficients of the sample-optimal boundary $\gamma$ and points on it, error matrices and $d^{\prime}_{b}$ values of classifying the samples using $\beta$ and $\gamma$, and the mapped scalar decision variables $\beta(\bm{x})$ and $\gamma(\bm{x})$ from the samples. The ability of scaling the $d^{\prime}$, by either warping the multivariate distributions, or their scalar decision variables, is available in this function in the options d_scale and d_scale_type. The function classify_normals_multi for classifying multiple normals.
Many different example problems, including every problem discussed in this paper (examples in figs. 2, 3, 4 and 5, tests in figs. 7 and 8, and research applications in fig. 9) are available as interactive demos in the ‘getting started’ live script of the toolbox, and can be easily adapted to other problems.
@@ 5 Performance benchmarks
In this section we test the performance (accuracy and speed) of our Matlab implementations of the generalized chi-square and ray-trace algorithms, against a standard Monte Carlo integration algorithm. We first set up a case where we know the ground truth, and compare the estimates of all three methods at the limit of high discriminability (low error rates) where it is most challenging (which occurs for computational models and ideal observers). We take two 3d normals with the same covariance matrix, so that the true discriminability $d^{\prime}$ is exactly calculated as their Mahalanobis distance. Now we increase their separation, while computing the optimal error with each of our methods at maximum precision, and a discriminability from it. The generalized chi-square method is very fast (due to the trivial planar boundary here), and the ray-trace method takes an average of 40s. For a fair comparison, we use $10^{8}$ samples for the Monte Carlo, which also takes $\sim$40s. Each method returns an estimate $\hat{d}^{\prime}$. Fig. 8a shows the relative inaccuracies $\lvert\hat{d}^{\prime}-d^{\prime}\rvert/d^{\prime}$ as true $d^{\prime}$ increases. With increasing separation, the Monte Carlo method quickly becomes inaccurate, since the error rate, i.e. the probability content in the integration domain, becomes too small to be sampled. The method stops working beyond $d^{\prime}\approx 10$, where none of the $10^{8}$ samples fall in the error domain. In contrast, inaccuracies in our methods are extremely small, of the order of the double-precision machine epsilon $\epsilon$, demonstrating that the algorithms contain no systematic error besides machine imprecision (however, Matlab’s native integration methods may not always reach the desired precision for a problem). This is possible because a variety of techniques are built into our algorithms to preserve accuracy, such as holding tiny and large summands separate to prevent rounding, using symbolic instead of numerical calculations, and using accurate tail probabilities. The inaccuracies do not grow with increasing separation, until $d^{\prime}\approx 75$, which corresponds to the smallest error rate $p(e)$ representable in double-precision ( realmin = 2e-308), beyond which both methods return $p(e)=0$ and $d^{\prime}_{b}=\infty$. For 1d problems beyond this, we can use $d^{\prime}_{e}$ instead. In a later paper [8], we use symbolic computation and variable precision arithmetic in the ray method to extend the computation of $d^{\prime}$ beyond 75, to values of around 70,000.